# Preference for Verifiability

### v1

@@ Abstract
Decision makers may face situations in which they cannot observe the consequences that result from their actions. In such decisions, motivations other than the expected utility of consequences may play a role. The present paper axiomatically characterizes a decision model in which the decision maker cares about whether it can be ex post verified that a good consequence has been achieved. Preferences over acts uniquely characterize a set of events that the decision maker expects to be able to verify in case they occur. The decision maker chooses the act that maximizes the expected utility across verifiable events of the worst possible consequence that may have occurred.
For example, a firm choosing between different carbon emission reduction technologies may find some technologies to leave ex post more uncertainty about the level of emission reduction than other technologies. The firm may care about proving to its stakeholders that a certain amount of carbon reduction has been achieved and may employ privately obtained evidence to do so. It may choose in expectation less efficient technologies if the achieved carbon reduction is better verifiable using the expected future evidence.
Keywords: Verifiability, uncertainty, maxmin utility, greenwashing, principal-agent
@@ 1 Introduction
In real-life decision making, we frequently make choices in which the final consequences we care about are only partially observable to us. For example, when deciding about different methods of CO2 compensation or emission reduction, the crucial consequence, the amount of CO2 emitted to the atmosphere, is not directly observed by most decision makers. Thus, the consequence the decision maker cares about remains more or less uncertain even after making a choice. In this case, it is natural that decision makers perceive a tradeoff between achieving good consequences in expectation and being able to verify that good consequences have been obtained.
This paper provides an attempt at a decision theory with potentially unobservable consequences. Most models in decision theory implicitly assume that after making a decision, decision makers can observe the states and consequences of their actions. In decisions under uncertainty without objective probabilities, commonly only ex ante uncertainty about consequences is permitted; after an act is chosen, there is no ex post uncertainty about the resulting consequence. While we employ the standard primitive of decision theory, preferences over acts, we dispose of the implicit assumption that there is no ex post uncertainty about the outcome. Removing this assumption and assuming decision makers care about ex post verifiability makes deviations from expected utility maximization plausible that can be naturally understood in the language of models from the ambiguity literature [[, e.g.,]]chandrasekher_dual_2022, gilboa_maxmin_1989, schmeidler_subjective_1989.
In particular, the dual-self interpretation of ambiguity aversion by [24] applies in our model, though with a different motivation and special functional form. Our model can be interpreted as an interaction between an expected utility maximizing agent –the decision maker– and a hidden principal, who’s actions are unobserved to the analyst. The principal and the decision maker have contrarian interests: one is trying to prove that the achieved outcome is good, the other that the achieved outcome is bad. It is this hidden interaction with the principal that can make the behavior of the decision maker appear ambiguity averse or ambiguity seeking.
The interaction between the decision maker and the hidden stakeholder unfolds like this: First, the decision maker chooses an act that maps states into consequences. Second, either the stakeholder or the decision maker receive various proofs about which state of the world may have obtained. A proof comes in the form of an event; a proof of a particular event is only obtained if the true state is indeed within the event. Third, whoever received the proofs chooses from the available proofs and confronts the other party with this evidence. Lastly, the one confronted with the evidence points to a state consistent with the proven event (and that is in their interest) and claims that this consequence is likely to have realized. To infer this hidden interaction, only preferences of the decision maker (over acts) have to be observed.
More concretely, consider the following example from corporate carbon responsibility (CCR). A firm chooses between different carbon emission offsetting programs. The efficacy of the different offsetting programs and thus the reduction of CO2 in the atmosphere is uncertain. Therefore, a carbon emission offsetting program is an act that maps uncertain states of the world (efficacy of methods, costs, etc.) into consequences (the amount of CO2 emitted to the atmosphere). Crucially, the exact state of the world (and thus the consequence) may remain unknown after the choice is made. The absence of objective probabilities and presence of ex post uncertainty may lead to greenwashing behavior: in their communication with stakeholders, firms may exaggerate the likelihood of states of the world in which their chosen action led to a large reduction in CO2. Alternatively, pressure from stakeholders may lead other firms to choose overly cautious CO2 offsetting programs because they want to be able to unambiguously prove that they have offset a certain amount of CO2. Both behaviors may lead to deviations from subjective expected utility maximization (and thus, welfare losses). Both behaviors also depend on the degree of ex post uncertainty; if additional information will become available pertinent to some actions but not others, firms may deliberately choose acts for which the information is more or less relevant. Firms potentially choose ex-ante inferior programs depending on whether for this program ex post more information will be available about the achieved emission reduction.
The present paper axiomatically characterizes such behavior. The decision maker is an expected utility maximizer over events that they expect to be able to verify. We call these events verifiable events and the set of verifiable events is subjective. For all other events, the decision maker exhibits extreme ambiguity aversion (or liking) and evaluates such events by the worst (or best) possible consequence that may have resulted. The case of extreme ambiguity aversion is interpreted as a preference for verification that good outcomes have been achieved. The case of extreme ambiguity liking is interpreted as a preference for obfuscation.
From the decision theoretic results we can derive policy implications in the context of CCR. Policy makers increasingly rely on CCR while focusing on regulating unverified environmental claims: for example, [28] plan to “proscribe […] generic environmental claims […] without proof” and “claims based on emissions offsetting schemes.” The comparative statics and welfare analysis of our models show that both greenwashing behavior as well as verification seeking behavior may lead to welfare losses and these losses are nonmonotone in the degree of lack of verifiability. In the principal agent relationshipe of CCR, increased transparency requirements only unambiguously improve welfare if full ex post verifiability is induced. If some ex post uncertainty remains, increased transparency may lead to a lower welfare, depending on the set of available actions of firms.
@@
The paper proceeds as follows: Section 2 introduces the example application of carbon offsetting programs in more detail. Section 3 introduces the notation before Section 4 introduces the two decision models. The axioms that characterize these models are introduced in Section 5. Comparative statics for the models can be found in Section 6. The paper also contributes a decision theoretic foundation to a model of greenwashing and on verifiability in contract theory. Such relations to the literature are discussed in Section 7.
@@ 2 Introductory Example
This section introduces an example of how the model presented in this paper can capture certain stylized facts about how firms interact with their stakeholders in corporate social responsibility issues where consequences are imperfectly observable. In such settings, subjective beliefs (for example about the efficacy of the firm’s actions) may greatly differ between stakeholders and firms and there might not exist an objective truth to relate to.
Suppose a firm decides between different carbon offsetting programs. One option is to purchase certificates of nature-based carbon capture, or more poignant, to plant Trees. Another option is to purchase renewable energy certificates, RECs, and the third option is to improve the production processes to increase Efficiency. For simplicity, we assume that the total budget to be spent on the offsetting programs is fixed.
@@
A greenwashing firm may choose to purchase RECs and may exaggerate in their communication to stakeholders the likelihood that state $t$ obtains. For example, [25] claim that for most firms joining the U.N. race to zero campaign [41], the CO2 emission reductions are not sufficiently verified. Firms may choose to purchase RECs even if they subjectively believe state $t$ to be very unlikely. Such behavior can be detected by an analyst from preferences that treat uncertain acts indifferent to their best possible consequence.
A verification seeking firm may choose Efficiency to eliminate any ex-post doubt of their stakeholders that they have achieved a certain amount of CO2 emission reductions. For example, a firm may try to prove to their stakeholders that their products are carbon neutral. Such firms may choose Efficiency even if they believe state $u$ to be very unlikely because ex post they cannot exclude this state. Such behavior can be detected from preferences that treat uncertain acts indifferent to their worst possible consequence.
In the following, we provide two models –and axiomatic characterizations thereof– that correspond to these two types of behavior. Compared to the extreme cases discussed above, allow for some events, for example $\{s,t\}$, to be ex-post verifiable and subjectively determined. From an applied perspective, the main contribution is to provide conditions on preferences over emission reduction methods from which greenwashing behavior can be distinguished from verification-seeking or expected utility behavior. In addition, for any firm following either preference, we can identify what information they expect to become available in the future.
@@ 3 Notation
Let $\mathscr{X}$ be a set of consequences and $\mathscr{S}$ a finite set of states of the world. $\mathscr{E}\subseteq 2^{\mathscr{S}}$ denotes the sigma algebra of events. A simple act is a measurable function $a:\mathscr{S}\rightarrow\mathscr{X}$ with a finite image. If $E$ is an event and $a,b$ are acts, then $a_{E}b$ is the act that agrees with $a$ on all states $s\in E$ and that agrees with $b$ on all states $s\in\overline{E}\equiv\mathscr{S}-E$. If $x\in X$, then $x$ also denotes the constant act that implements $x$ in all states.
@@
An event $E$ is said to be null if $\beta\sim\gamma E\beta$ for all $\gamma,\beta\in X$ such that $\gamma\succ\beta$. An event $E$ is said to be universal if $\gamma\sim\gamma E\beta$ for all $\gamma,\beta\in X$ such that $\gamma\succ\beta$. An event $E$ is said to be essential if $\gamma\succ\gamma E\beta\succ\beta$ for some $\gamma,\beta\in\mathscr{X}$.
Unlike in expected utility, in our axiomatization what happens on null events may in principle still be preference relevant. We therefore introduce the stronger notion of irrelevant events. An event $E$ is said to be irrelevant if $\gamma_{E}a\sim a$ for all $\gamma\in\mathscr{X}$ and $a\in\mathscr{A}$. A state of the world is said to be irrelevant if it is an element of an irrelevant event. The set of events that are not irrelevant is denoted $\mathscr{E}^{*}$ and the set of states that are not irrelevant is denoted $\mathscr{S}^{*}$. An event that is null need not be irrelevant. In our model, only if an event is irrelevant, then it never matters for preferences what consequences are acquired on this event.
@@ 4 Decision Models
We analyze two decision models. In the first decision model, the decision maker has the desire to verify that a particular utility level has been reached. In the context of our example, this may arise because the decision maker wants to prove to a stakeholder that a certain benefit of taking the action has been achieved.
Definition 1 (Expected Verification Utility).
@@
$\mathscr{V}$ is called the set of verifiable events. If $\mathscr{V}=\mathscr{E}$, then every state is verifiable and the decision maker maximizes expected utility. The verifiable events do not necessarily form a partition but are closed under intersections. It follows that the set of verifiable events is a $\pi$-system that contains $\mathscr{S}^{*}$. The interpretation of a verifiable event $E$ is that if the event obtains (i.e., if the true state of the world is within $E$), then the decision maker receives a proof that they can use to prove that this event obtained. They then use this proof to prove to the stakeholder that at least the consequence $\min_{t\in E}u(a(t))$ has been achieved. [^1] The model assumes the verifiable events to be act-independent. The act-dependent case is not treated in the current paper for two reasons: first, if such verifiable events would be objectively given for every act, then this would make the analysis almost trivial. Second, if there are subjectively determined act-dependent verifiable events, then almost any behavior is possible and the verifiable events are unlikely to have interesting uniqueness properties.
@@
Firm A would like to be able to prove to its stakeholders that a high level of reduction of CO2 emissions has been achieved. The firm believes that state $s$ and $u$ are unlikely, though not impossible. In case the event $\{s,t\}$ obtains, the firm believes it is able to prove this to its stakeholders. For example, after purchasing any amount of certificates of any kind, the firm can show these certificates to its stakeholders. However, the firm is unable to prove that $\{u\}$ obtains if certificate prices are not publicly observable – it cannot use a low number of purchased certificates to prove that it is impossible to buy more certificates. The firm is also unable to prove that $\{s\}$ or that $\{t\}$ obtains since proving this would require observing the hypothetical behavior of how other market participants would behave given a purchase of RECs or no purchase of RECs. The firm can however verify that $\{s,t,u\}$ obtains since there is no other nonnull state. Firm A therefore considers two events when making a decision, $\{s,t\}$ and $\{s,t,u\}$. With probability $\mu(\{u\})$ it can only prove that $\{s,t,u\}$ obtains and that the emission reduction is at least equal to the worst consequence of an act. With probability $\mu(\{s,t\})$ the firm can verify that $\{s,t\}$ obtains and that the emission reduction is at least equal to the worst possible consequence on states $s$ and $t$. The firm therefore multiplies the probability of $u$ with the utility of the worst possible consequence on $\{s,t,u\}$ and the probability of $\{s,t\}$ with the worst possible consequence on $\{s,t\}$. If the probability of state $u$ is sufficiently low, the firm will choose Trees. If $u$ is sufficiently likely, the firm will choose Efficiency. ∎
In the second decision model, the decision maker has the desire to obfuscate that a bad consequence might have resulted. In the context of our example, this may arise because a stakeholder may confront the firm with evidence about the state of the world (e.g., that $\{s\}$ obtains after having chosen RECs).
@@
The interpretation of this model is that after a stakeholder confronts the decision maker with the proof that an event $E$ obtains, the decision maker points to the best possible consequence that might have been achieved on $E$. We assume that the stakeholder has adversarial preferences to the decision maker, i.e., the stakeholder tries to prove that at most a certain amount of utility has been obtained. The stakeholder always uses all available verifiable information (chooses the smallest event in $\mathscr{V}$ that contains the true state) and the decision maker then always points to the best consequence on this event.
@@
Both of these decision models are extreme cases and unobservable consequences may generate a rich variety of decision models that are less extreme. The present decision models only represent a starting point for the exploration of decision theories in which consequences cannot be observed.

### v2

# arXiv:2508.19585v2[econ.TH] 02 Mar 2026

@@ Abstract
Decision makers sometimes cannot observe the consequences of their actions ex-post. This paper axiomatically characterizes a decision model in which the decision maker cares about verifying that a good consequence has been achieved. Preferences over acts identify a set of events the decision maker expects to verify. Decision makers choose acts maximizing, in expectation over verifiable events, the worst-case utility consistent with each event. A dual model captures decision makers who instead seek to obscure poor outcomes from verification. As an application, firms choosing carbon-reduction technologies may prefer less efficient but more verifiable technologies to prove emission reductions to stakeholders.
Keywords: Verifiability, uncertainty, Choquet expected utility, greenwashing, principal-agent
@@ 1 Introduction
In many real-world decisions, the final consequences are only partially observable. For instance, when a firm chooses a method for CO2 emission reduction, the actual amount of CO2 removed from the atmosphere is not directly observed but must be inferred from indirect evidence, such as carbon compensation certificates. This lingering ex-post uncertainty about consequences can motivate choices that deviate from standard expected utility theory. A firm might not simply aim to maximize the expected reduction in emissions; its incentives are often more complex due to the presence of stakeholders. The firm may want to prove that a significant emission reduction has been achieved, or conversely, it may want to avert blame if stakeholders can prove a desirable target was missed.
This paper’s central theoretical contribution is to model these motivations by providing an axiomatic foundation for preferences that depend on the ex-post verifiability of outcomes. We ask: how does the anticipation of what can be proven after the fact shape choice under uncertainty? We show that behavior driven by a preference for or against ex-post verifiability can be naturally understood through the lens of ambiguity aversion models fulfilling comonotonic independence [undefap], which have traditionally been motivated by ex-ante uncertainty. We rationalize ambiguity attitudes as a coherent response to unobservable consequences.
The applications we consider, such as corporate environmental claims, are characterized by a lack of objective probabilities. This subjective environment often gives rise to phenomena like “greenwashing,” where claims about expected emission reductions greatly differ between firms and stakeholders. Consequently, we adopt a fully subjective Savage framework with subjective mixtures [undefae] to identify risk preferences.
We propose two models that capture the principal-agent relationships that can arise from imperfect ex-post verifiability using a dual-self representation similar to [undefy]. The first model, certification utility, describes a decision-maker who wants to prove a good outcome was achieved. They evaluate acts based on the expected utility of the outcome they can prove, i.e., the worst outcome consistent with their best proof. This captures the behavior of a firm genuinely seeking to demonstrate its positive environmental impact and corresponds to extreme ambiguity aversion. The second model, obfuscation utility, captures the opposite motive. Here, a stakeholder presents evidence, and the firm tries to point to the most favorable interpretation consistent with that evidence. This model provides a behavioral foundation for greenwashing, where firms exploit ambiguity to avoid blame for poor outcomes, and corresponds to extreme ambiguity seeking.
Our main representation theorems show that from the decision maker’s choices alone, an outside observer can identify the key elements of the model: the presence of a hidden stakeholder, the nature of their ex-post interaction (verification or obfuscation), and the set of events the decision maker anticipates being verifiable along with the probabilities of these events. This means we can distinguish between a firm that is an expected utility maximizer, one that is genuinely seeking verifiability, and one that is engaging in obfuscation, simply by observing their preferences. The endogenously derived set of verifiable events aligns with standard assumptions in contract theory, being closed under intersections but not necessarily under complements.
Our decision-theoretic results have policy implications for corporate carbon responsibility (CCR). As policymakers seek to regulate unverified environmental claims—for instance, by planning to “proscribe […] generic environmental claims […] without proof” and “claims based on emissions offsetting schemes” [undefac]—our framework provides a tool to analyze the behavioral consequences. A welfare analysis reveals that both certification-seeking and obfuscating behaviors can lead to welfare losses. These losses are non-monotonic in the degree of verifiability; that is, increasing transparency requirements may lower welfare unless full ex-post verifiability is achieved. This suggests that policies promoting CCR and transparency must be designed with care, as their effectiveness depends critically on the set of actions available to firms.
@@
The paper proceeds as follows: Section 2 introduces the example application of carbon offsetting programs in more detail. Section 3 presents the notation before Section 4 discusses the two decision models. The axioms that characterize these models are developed in Section 5. Comparative statics for the models can be found in Section 6. The paper also contributes a decision-theoretic foundation to a model of greenwashing and on verifiability in contract theory. Such relations to the literature are discussed in Section 7.
@@ 2 Introductory Example
This section introduces an example of how the model presented in this paper can capture certain stylized facts [^1] In the U.N. Race to Zero Campaign [undefar], an analysis of the largest firms of eight sectors [undefaj] found that most firms relied on renewable energy certificates to offset carbon emissions while only a small number of firms used certificates of nature-based carbon capture. [undefaj] concluded that “Companies’ climate pledges for 2030 fall well short of the required ambition and are inappropriately verified” (p. 6), that “the climate strategies of 15 of the 24 companies [are] of low or very low integrity” and that “targets and potential offsetting plans remain ambiguous” (p. 5). about how firms interact with their stakeholders in CCR issues where beliefs are subjective and consequences are imperfectly observable, such as carbon emission reductions. In such settings, subjective beliefs (for example about the efficacy of the firm’s actions) may greatly differ between stakeholders and firms and there usually does not exist an objective probability distribution to relate to. If firms and stakeholders cannot agree ex-ante what the best course of action is, this leaves only the ex-post outcomes for the stakeholder to evaluate whether the firm has done well.
Suppose a firm decides between different carbon offsetting programs. One option is to purchase certificates of nature-based carbon capture, or more pointedly, to plant Trees. Another option is to purchase renewable energy certificates, RECs, and the third option is to improve the production processes to increase Efficiency. For simplicity, we assume that the total budget to be spent on the offsetting programs is fixed.
@@
A greenwashing firm may choose to purchase RECs and may exaggerate in their communication to stakeholders the likelihood that state $t$ obtains. For example, [undefaj] claim that for most firms joining the U.N. race to zero campaign [undefar], the CO2 emission reductions are not sufficiently verified. Firms may choose to purchase RECs even if they subjectively believe state $t$ to be very unlikely. Such behavior can be detected by an analyst from preferences under which uncertain acts are indifferent to their best possible consequence.
A certification-seeking firm may choose Efficiency to eliminate any ex-post doubt of their stakeholders that they have achieved a certain amount of CO2 emission reductions. For example, a firm may try to prove to its stakeholders that its products are carbon neutral. Such firms may choose Efficiency even if they believe state $u$ to be very unlikely because ex post they cannot exclude this state. If however the event $\{s,t\}$ would be ex-post verifiable, they may choose Trees instead. Such behavior can be detected from preferences that treat uncertain acts indifferent to their worst possible consequence.
In the following, we provide two models —and axiomatic characterizations thereof— that correspond to these two types of behavior. Compared to the extreme cases discussed above, we allow for some events, for example $\{s,t\}$, to be ex-post verifiable and subjectively determined. From an applied perspective, the main contribution is to provide conditions on preferences over emission reduction methods from which greenwashing behavior can be distinguished from certification-seeking or expected utility behavior. In addition, for any firm following either preference, we can identify what information they expect to become available in the future.
@@ 3 Notation
Let $\mathscr{X}$ be a set of consequences and $\mathscr{S}$ a finite set of states of the world. $\mathscr{E}\subseteq 2^{\mathscr{S}}$ denotes the sigma algebra of events. A simple act is a measurable function $a:\mathscr{S}\rightarrow\mathscr{X}$ with a finite image. If $E$ is an event and $a,b$ are acts, then $a_{E}b$ is the act that agrees with $a$ on all states $s\in E$ and that agrees with $b$ on all states $s\in\overline{E}\equiv\mathscr{S}-E$. If $x\in\mathscr{X}$, then $x$ also denotes the constant act that implements $x$ in all states.
@@
An event $E$ is said to be null if $\beta\sim\gamma E\beta$ for all $\gamma,\beta\in\mathscr{X}$ such that $\gamma\succ\beta$. An event $E$ is said to be universal if $\gamma\sim\gamma E\beta$ for all $\gamma,\beta\in\mathscr{X}$ such that $\gamma\succ\beta$. An event $E$ is said to be essential if $\gamma\succ\gamma E\beta\succ\beta$ for some $\gamma,\beta\in\mathscr{X}$.
Unlike in expected utility, in our axiomatization what happens on null events may in principle still be preference relevant. We therefore introduce the stronger notion of irrelevant events. An event $E$ is said to be irrelevant if $\gamma_{E}a\sim a$ for all $\gamma\in\mathscr{X}$ and $a\in\mathscr{A}$. A state of the world is said to be irrelevant if it is an element of an irrelevant event. The set of events that are not irrelevant is denoted $\mathscr{E}^{*}$ and the set of states that are not irrelevant is denoted $\mathscr{S}^{*}$. An event that is null need not be irrelevant. In our model, an event is only irrelevant for the decision maker’s preference if it is irrelevant in this stronger sense we defined here.
@@ 4 Decision Models
We analyze two decision models. In the first decision model, the decision maker wants to verify that a particular utility level has been reached. In the context of our example, this may arise because the decision maker wants to prove to a stakeholder that a certain benefit of taking the action has been achieved.
Definition 1 (Expected Certification Utility).
@@
$\mathscr{V}$ is called the set of verifiable events. If $\mathscr{V}=\mathscr{E}$, then every state is verifiable and the decision maker maximizes expected utility. The verifiable events do not necessarily form a partition but are closed under intersections and thus form a $\pi$-system that contains $\mathscr{S}^{*}$. The interpretation of a verifiable event $E$ is that if the event obtains (i.e., if the true state of the world is within $E$), then the decision maker receives a proof that they can use to prove that this event obtained. They then use this proof to convince the stakeholder that at least the consequence $\min_{t\in E}u(a(t))$ has been achieved. [^2] The model assumes the verifiable events to be act-independent. The act-dependent case is not treated in the current paper for two reasons: first, if such verifiable events would be objectively given for every act, then this would make the analysis almost trivial. Second, if there are subjectively determined act-dependent verifiable events, then almost any behavior is possible and the verifiable events are unlikely to have interesting uniqueness properties.
@@
Firm A would like to be able to prove to its stakeholders that a high level of reduction of CO2 emissions has been achieved. The firm believes that states $s$ and $u$ are unlikely, though not impossible. In case the event $\{s,t\}$ obtains, the firm believes it is able to prove this to its stakeholders. For example, after purchasing any amount of certificates of any kind, the firm can show these certificates to its stakeholders. However, the firm is unable to prove that $\{u\}$ obtains if certificate prices are not publicly observable – it cannot use a low number of purchased certificates to prove that it is impossible to buy more certificates. The firm is also unable to prove that $\{s\}$ or that $\{t\}$ obtains since proving this would require observing the hypothetical behavior of how other market participants would behave given a purchase of RECs or no purchase of RECs. The firm can however verify that $\{s,t,u\}$ obtains since there is no other nonnull state. Firm A therefore considers two events when making a decision, $\{s,t\}$ and $\{s,t,u\}$. With probability $\mu(\{u\})$ it can only prove that $\{s,t,u\}$ obtains and that the emission reduction is at least equal to the worst consequence of an act. With probability $\mu(\{s,t\})$ the firm can verify that $\{s,t\}$ obtains and that the emission reduction is at least equal to the worst possible consequence on states $s$ and $t$. The firm therefore multiplies the probability of $u$ with the utility of the worst possible consequence on $\{s,t,u\}$ and the probability of $\{s,t\}$ with the worst possible consequence on $\{s,t\}$. If the probability of state $u$ is sufficiently low, the firm will choose Trees. If $u$ is sufficiently likely, the firm will choose Efficiency. ∎
In the second decision model, the decision maker has the desire to obfuscate whether a bad consequence might have resulted. In the context of our example, this may arise because a stakeholder may confront the firm with evidence about the state of the world (e.g., that $\{s\}$ obtains after having chosen RECs).
@@
The interpretation of this model is that after a stakeholder confronts the decision maker with the proof that an event $E$ obtains, the decision maker points to the best possible consequence that might have been achieved on $E$. We assume that the stakeholder has adversarial preferences toward the decision maker, i.e., the stakeholder tries to prove that at most a certain amount of utility has been obtained. The stakeholder always uses all available verifiable information (chooses the smallest event in $\mathscr{V}$ that contains the true state) and the decision maker then always points to the best consequence on this event. This model is dual to the certification utility in the sense that for a certification utility minimizing decision maker there exists an expected obfuscation utility with a reverse preference over outcomes and vice versa.
@@
Both of these decision models are extreme cases and unobservable consequences may generate a rich variety of decision models that are less extreme. The present decision models only represent a starting point for the exploration of decision theories in which consequences cannot be observed. The following axioms are therefore to be understood as a starting point for a systematic analysis of preferences under uncertainty with imperfect ex-post verifiability of consequences.